OpenStudy (amour.anohs):

Henry is 3 times older than Julia. Rhianna is 1/2 as old as Julia. What formula shows the relationship between Henry and Rhianna's ages?

OpenStudy (amour.anohs):

Henry is 3 times older than Julia: H = 3J J = H/3 Rhianna is 1/2 as old as Julia: R = 1/2J J = 2R Now, set the equations equal to each other: H/3 = 2R Answer:H = 6R

candycove (candycove):

omg